the sum of three numbers is 254. the second is 3 times the first, and the third is 5 less than the second. find the numbers.
Answer by checkley77(12844)

second=3f
third=s-5
f=first
f+3f+(3f-5)=254
7f-5=254
7f=254+5
7f=259
f=259/7
f=37 is the first number.
s=3*37=111 is the second number.
t=111-5=106 is the third number.
Proof:
37+111+106=254
254=254